OpenStudy (anonymous):

Can anyone help with some factoring ? /:

OpenStudy (anonymous):

I don't need the answer but I need someone to help me guide through this , its confusing me /:

OpenStudy (anonymous):

it is confusing to me too, all this verbiage. but we can try "factor by grouping" for \[2x^2 + 13x + 15 \]

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

we will break the middle term in to two pieces lets try \(12x+x\) if it works great, if not we will change it

OpenStudy (anonymous):

yeah i'll help you with some factoring ;)

OpenStudy (anonymous):

wait what did u just do ?

OpenStudy (anonymous):

\[2x^2+12x+x+15\] is our first start then from \(2x+12x\) we can factor out a \(2x\) to get \[2x(x+6)+x+15\] but you see that the second possible factor is \(x+15\) which is not equal to \(x+6\) so lets try again

OpenStudy (anonymous):

i looked at \{13x\) and broke it in to two parts as \(13x=12x+x\)

OpenStudy (anonymous):

but that didn't work, often the first guess is wrong

OpenStudy (anonymous):

lets try \(13x=10x+3x\)

OpenStudy (anonymous):

then we write \[2x^2+10x+3x+15\] and factor \(2x\) out of the first two terms, that is, factor \(2x\) out of \(2x^2+10x\)

OpenStudy (anonymous):

you get \[2x^2+10x+3x+15=2x(x+5)+3x+15\] are you okay to this step?

OpenStudy (anonymous):

you don't have to solve it , you just have to explain how a trinomial of the form 2x2 + 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping.

OpenStudy (anonymous):

for part 1 , and then part 2 we can make up our own problem .

OpenStudy (anonymous):

you can turn it in two a four term polynomial by rewriting it as \[2x^2+10x+3x+15\]

OpenStudy (anonymous):

that is because \(2x^2+10x\) has a common factor of \(2x^2\) and \(3x+15\) has a common factor of \(3\)

OpenStudy (anonymous):

ohh alright now I see it

OpenStudy (anonymous):

so you can factor as \[2x^2+13x+15=2x^2+10x+3x+15=2x(x+5)+3(x+5)\]\[=(2x^2+3)(x+5)\]

OpenStudy (anonymous):

oops typo there, last line should be \[(2x+3)(x+5)\] sorry

OpenStudy (anonymous):

its fine haha :] thanks for your help wanna help me with part 2?

OpenStudy (anonymous):

to answer part 2, you can make up your own polynomial that factor by first writing it in factored form say you pick \((3x+1)(x+2)\)

OpenStudy (anonymous):

then when you multiply it out you would get \[3x^2+6x+x+2=3x^2+7x+2\]and since you started with the factored form you know that \[3x^2+7x+2\] will factor

OpenStudy (anonymous):

when you multiplied that you foiled it right ? im just trying to make sure I understand what you did there .